OpenStudy (anonymous):

Find the equation of the line passing through the points (-2, 5) and (3, 10).

OpenStudy (anonymous):

y = x +7

OpenStudy (anonymous):

y - y1 = m(x - x1) m = (y2-y1)/(x2-x1) x1=-2,y1=5,x2=3,y2=10 y - 5 = (10-5)(x+2)/(3+2) y - 5 = x + 2 y = x + 2 + 5 y = x + 7

OpenStudy (anonymous):

Thankyou so much. Very nice layout .